# Reduction of Forgetting by Contextual Variation During Encoding Using 360-Degree Video-Based Immersive Virtual Environments

### v1

@@ I Introduction
We experimented based on the experimental design adopted by Smith and Handy [5, 14], which allowed us to examine the reinstatement and decontextualization effects simultaneously. The participants studied 20 Tagalog-Japanese pairs, performed five successive retrieval practices, and completed a final cued recall test two days later. We primarily compared the amount of forgetting from the first day (Day 1) to the second day (Day 2) and hypothesized that learning in the same IVE on Day 1 as that on Day 2 would reduce forgetting due to the reinstatement effect and that learning in various IVEs on Day 1 would prevent forgetting due to the decontextualization effect.
@@ II-B Decontextualization
To explain the decontextualization, two hypotheses have been proposed. One hypothesis is encoding variability [23], which states that an increase in the number of contextual cues attached to the memory to be remembered facilitates recall. For example, in the experiment by Smith et al. [4], there were twice as many contextual cues supporting retrieval under the two-room condition as under the one-room condition. The other hypothesis is desirable difficulty [24], which states that more information is encoded while retrieval is attempted when a recall is inhibited. In the varied condition, relearning occurs in contexts different from the original context, a loaded condition in which context-dependent forgetting is likely to occur, such that more information is stored in the memory.
@@ II-C Environmental Context Manipulation Using Virtual Reality
Additionally, previous studies on context-dependent effects using VR have not fully examined important indicators of the VR experience. The most important measure is presence. Presence is the sense of being in an IVE [15]. The degree of presence is known to affect several aspects of cognition and behavior in IVE. Specifically, Smith reviewed studies of episodic memory using VR and suggested the close relationship between episodic memory and presence [29]. This relationship may be based on the fact that presence and episodic memory are related to spatial recognition and depend on the functioning of the hippocampus [13, 30]. Context-dependent effects are one of the phenomena of episodic memory, which is likely to be affected by presence. However, this has not been investigated in previous studies. Therefore, this study measures presence using questionnaires and discusses its relationship with context-dependent effects.
@@ III-E Environmental Contexts
Figure 4: Shots of 360-degree videos used as environmental context. From top left: Beach, Pond, River, Shrine, Park, Building, and Glassland. The shots were taken by a KANDAO Qoocam8K camera with a resolution of 7680 $\times$ 3840, 30 fps, and monoscopic.
@@ V-D Forgetting
Figure 6: Forgetting from Day 1 to Day 2. Error bars represent the 95% confidence intervals. (a) the number of forgotten words. (b) the ratio of forgotten words. For both measures, forgetting was significantly lower in the Varied condition than in the other two conditions, confirming the decontextualization effect. $*{}^{*}$: p $<$ .05.
@@ V-E Presence
During the adaptation section of Day 1, the participants answered IPQs once for each of the seven IVEs. List-wise, we removed the data of seven participants owing to missing information. The IPQ scores, or presence, for the seven IVEs are presented in Fig. 7a. A Friedmann test showed no significant difference between the IVEs ($\chi^{2}$ (6) = 6.61, p = .36).
@@ V-F VR Sickness
After conducting the experiment for two days, the participants answered SSQ to indicate the degree of VR sickness felt by them. List-wise, two participants were removed and were not included in the analysis because they were missing some values. The SSQ total scores (SSQ-TS) obtained on Day 1 and Day 2 in each context condition are presented in Fig. 9. After aligned rank transformation [40], three-by-two mixed ANOVA was computed, with context (Constant-NotReinstated, Constant-Reinstated, and Varied) as a between-participants variable and measurement timing (Day 1 and Day 2) as a repeated measure. The result showed that only the main effect of measurement timing was significant (F (1, 58) = 132.9, p $<$ .001, ${\eta_{p}}^{2}$ = 0.70). Conversely, the main effect of context and the interaction were not significant (F (2, 58) = 1.09, p = .34, ${\eta_{p}}^{2}$ = 0.04; F (2, 58) = 0.37, p = .69, ${\eta_{p}}^{2}$ = 0.01).
@@ VI-A No reinstatement effect using IVE context
In this study, we investigated the reinstatement effect on all memory measures (RP1, acquisition, retention, and forgetting) and found no significant effects. Therefore, the hypotheses were not supported. Whereas the previous studies used 3DCG IVEs [9, 7, 6], this study is the first to use live-action 360-degree video for IVE. We assumed that increased realism and familiarity improved the effectiveness of the environmental context, but the assumption was not supported.
@@
For RP1 and acquisition, the hypothesis stated that the number of correct responses obtained would be higher in the same context conditions (Constant-NotReinstated and Constant-Reinstated conditions), where tests were conducted in the same context as the initial study than in the different context condition (Varied condition), where tests were conducted in a new context. From Fig. 5, we observed no differences between the conditions, which were very different from those observed by Smith and Handy [5, 14] who showed the superiority of the same context condition to the different context condition by video/photo context manipulation in a similar procedure. Particularly in RP1, the number of correct answers tended to be slightly higher in the different context condition than in the same context condition (same context: M = 3.71, SD = 2.52; different context: M = 4.05, SD = 3.01; Cohen’s d = 0.12), which was the opposite to the hypothesis. Thus, the result of no reinstatement effect observation was not obtained due to insufficient detection power.
@@
For retention, the hypothesis stated that the number of correct responses would be higher in the Constant-Reinstated condition, where participants performed the final test in the same context as Day 1, than in the Constant-NotReinstated condition, where the participants performed the final test in a context different from that of Day 1. Similarly, for forgetting, the hypothesis stated that forgetting would be lower in the Constant-Reinstated condition than in the Constant-NotReinstated condition. From Fig. 5 and Fig. 6, we observed no significant differences between the conditions, and the hypothesis was not supported. However, unlike in the case of RP1 and acquisition, the Constant-Reinstated condition was numerically dominant. Although the effect size was small, the detection power may have been insufficient.
@@
The lack of effect may have been caused by cue overload [41], which is a common phenomenon observed in this field of study. Although the use of too many words eliminates the reinstatement effect, no such occurrence was analyzed in our experiment because Smith and Handy [14], who also used 20 Tagalog-English pairs of words, detected the effect. On the other hand, some procedural differences were observed between our study and that of Smith and Handy [14]. For example, Smith and Handy quickly switched the background photo contexts (one photo/word) in their study, while we used the same context for 20 items. The experimental protocol was modified because quick changes in IVE could confuse the participants and compromise the sense of presence, which is vital for inducing context-dependent effects using VR. This difference in cue load level was not problematic in that it avoided the problem of cue overload but may have reduced the reinstatement effect [12]. However, Wälti et al. [6] already proved that even in a method like Smith and Handy’s, where the contexts are changed quickly per word, no reinstatement effect can be obtained in context manipulation by VR. Therefore, we do not consider the effect to cause a severe problem.
@@
Instead, the unique features of VR may have weakened the effect. From the discussion presented by Watson and Gaudl [8], we deduced that the congruence of the meta-cognitive context of “experiencing VR” may have counteracted the effects of visual and auditory context changes. This “experiencing VR” context was thought to be formed by wearing the HMD and high arousal during the VR experience. Considering the assumption to be correct, the context cannot be changed irrespective of the extent of audio-visual environmental information manipulation in VR. Similarly, the congruence of the meta-cognitive context of “being at the same university and undergoing the same experiment” was suggested to weaken the context-dependent effect when using the room as the environmental context [27]. However, despite the meta-cognitive context match, ways in which the effect can be stably detected have been suggested, for example, by using the background video context [12] or simultaneous manipulation of place, task, and social context [42]. Moreover, as the decontextualization effect was confirmed in the present study, the meta-cognitive contextual congruence of experiencing VR is suggested not to be fatal. Therefore, further research is expected to reveal conditions under which the reinstatement effect occurs using VR despite the meta-cognitive context match.
@@
One method to solve the problem could employ a further increase in presence. In this study, the group of participants who reported higher presence tended to show a more evident reinstatement effect than those who reported lower presence, as shown in Fig. 8. Although this trend was not statistically significant, the fact that the prevailing conditions were switched is noteworthy. The high-presence group forgot less in the Constant-Reinstated condition, while the low-presence group forgot less in the Constant-NotReinstated condition. Increasing body tracking levels may be considered effective for enhancing presence [30]. As only three tracking points were used (the head and both hands) in this study, presence can be improved by attaching trackers to the waist and feet of the participants to achieve full-body tracking. Further, if the movements of the fingers can be faithfully reflected in the avatar by hand tracking, the presence can be improved. However, the possibility of the relationship between presence and context-dependent effect indicated that participants, who more likely felt the presence, were more likely to be context-dependent because both characteristics are dependent on hippocampal activity [13, 30]. A test to manipulate presence as an independent variable should be conducted in a future study, considering its importance.
@@ VI-B Decontextualization effect using IVE context
The second difference between our study and that of Watson and Gaudl [8] includes the type and number of IVEs used. Watson and Gaudl used two photorealistic 3DCG indoor environments, while we used seven 360-degree-video outdoor environments. Based on the encoding variability hypothesis with respect to the number of contexts, the more environmental contexts used, the more were cues available. This result was expected to improve the recall performance [43]. Conversely, the differences between 3DCG and 360-degree video, indoor and outdoor, can also be verified in the future.
Finally, Watson and Gaudl used doors for a contextual change, as doors have been found to act as boundaries for episodic memory [44]. Conversely, the participants of this study remained seated and only moved to put on and off their HMD. Therefore, the possibility that putting on and off the HMD acted as an episodic boundary similar to a door or visually representing movement (e.g., virtual doors or portals) in the IVE may have improved the effect remains to be explored.
The difference between the Constant-NotReinstated and Varied conditions tended to be larger for participants with higher presence, as shown in Fig. 8. Although this trend was not statistically significant, we consider it to be an essential hint for further improving the memory support effect obtained in this study. According to the aforementioned discussion on the reinstatement effect, we can manipulate presence as an independent variable for validation in the future; for instance, by manipulating the tracking level.
Finally, the VR sickness measured by the SSQ did not differ significantly between conditions, as shown in Fig. 9. Therefore, the characteristic was not confounded for memory performance. VR sickness is one of the critical factors to be considered during the practical application of VR-based systems. Therefore, this study is noteworthy in that we measured VR sickness for the first time in environmental context-dependent memory research using VR and showed relatively low severity.
@@ VI-C Limitations
Additionally, this study demonstrated a new method of foreign language learning as a clear application by employing the task of memorizing Japanese translations of Tagalog words. However, the generalizability of this research findings may go beyond language learning. For example, Smith and Handy (2014) showed decontextualization effects in a face/name pair memory task [5], and it is highly possible that our findings can be applied to various areas as long as classified in the field of episodic memory. On the other hand, the results are unlikely to be applicable to memories that cannot be described verbally, such as memories of motor skills. We employed language learning to be consistent with the use of word lists as the material to be memorized in most previous studies and also because Smith and Handy (2016) [14], who found decontextualization effects, employed Tagalog-English pair learning. Future studies are expected to conduct validation in various situational settings to clarify the generalizability of the results.
@@ VI-D Future Work
In this experiment, it was verified that while the reinstatement effect using VR is unlikely to occur, the decontextualization effect does occur. As the decontextualization effect using VR was revealed for the first time in this study, replication and follow-up experiments should be conducted to elucidate the conditions under which the effect can be enhanced. For example, the experiment in this study suggested that the context-dependent effect may become more apparent in participants with higher presence. The relationship between presence and context-dependent effect is, therefore, essential for future work.
@@ VII Conclusion
This study tested whether context-dependent forgetting can be eliminated by inducing reinstatement and decontextualization effects using 360-degree video-based virtual environments. For the reinstatement effect, complex results were obtained depending on the dependent variables, suggesting the need for further elaborate research. On the other hand, for the decontextualization effect, this study was the first to validate the effect using VR and revealed its statistically significant effect with a large effect size. While previous studies have assumed that VR could naturally reproduce the context-dependent effects observed in the real world, most of them could not prove the assumption. Given this context, obtaining a clear decontextualization effect in this study sheds light on the possibility of reproducing context-dependent effects using VR.

### v2

@@ I Introduction
We employed the experimental design adopted by Smith and Handy [5, 14], which allowed us to examine the reinstatement and decontextualization effects simultaneously. The participants studied 20 Tagalog-Japanese pairs, performed five successive retrieval practices, and completed a final cued recall test two days later. We primarily compared the amount of forgetting from the first day (Day 1) to the second day (Day 2) and hypothesized that learning in the same IVE on Day 1 as that on Day 2 would reduce forgetting due to the reinstatement effect and that learning in various IVEs on Day 1 would prevent forgetting due to the decontextualization effect.
@@ II-B Decontextualization
Two hypotheses have been proposed to explain the decontextualization. One hypothesis is encoding variability [23], which states that an increase in the number of contextual cues attached to the memory to be remembered facilitates recall. For example, in the experiment by Smith et al. [4], there were twice as many contextual cues supporting retrieval under the two-room condition as under the one-room condition. The other hypothesis is desirable difficulty [24], which states that more information is encoded while retrieval is attempted when a recall is inhibited. In the varied condition, relearning occurs in contexts different from the original context, a loaded condition in which context-dependent forgetting is likely to occur, such that more information is stored in the memory.
@@ II-C Environmental Context Manipulation Using Virtual Reality
Additionally, previous studies on context-dependent effects using VR have not fully examined important indicators of the VR experience. The most important measure is presence. Presence is the sense of being in an IVE [15]. The degree of presence is known to affect several aspects of cognition and behavior in IVE. Specifically, Smith reviewed studies of episodic memory using VR and suggested the close relationship between episodic memory and presence [29]. This relationship may be based on the fact that presence and episodic memory are related to spatial recognition and depend on the functioning of the hippocampus [13, 30]. Context-dependent effects are one of the phenomena of episodic memory, which is likely to be affected by presence. However, this has not been investigated in previous studies. Therefore, this study measures presence using questionnaires and discusses its relationship with context-dependent effects. Although we employed a widely used questionnaire to measure presence in VR research, it should be noted that there is a debate as to whether such a questionnaire can truly measure presence [31].
@@ III-E Environmental Contexts
Figure 4: Shots of 360-degree videos used as environmental context. From top left: Beach, Pond, River, Glassland, Shrine, Park, and Building. The shots were taken by a KANDAO Qoocam8K camera with a resolution of 7680 $\times$ 3840, 30 fps, and monoscopic.
@@ V-D Forgetting
Figure 6: Forgetting from Day 1 to Day 2. Error bars represent the 95% confidence intervals. (a) the number of forgotten words. (b) the ratio of forgotten words. For both measures, forgetting was significantly lower in the Varied condition than in the other two conditions, confirming the decontextualization effect. : p $<$ .05.
@@ V-E Presence
During the adaptation section of Day 1, the participants answered IPQs once for each of the seven IVEs. List-wise, we removed the data of seven participants owing to missing information. The IPQ scores for the seven IVEs are presented in Fig. 7a. A Friedmann test showed no significant difference between the IVEs ($\chi^{2}$ (6) = 6.61, p = .36).
@@ V-F VR Sickness
After conducting the experiment for two days, the participants answered SSQ to indicate the degree of VR sickness they felt. List-wise, two participants were removed and excluded from the analysis because they were missing some values. The SSQ total scores (SSQ-TS) obtained on Day 1 and Day 2 in each context condition are presented in Fig. 9. After aligned rank transformation [41], three-by-two mixed ANOVA was computed, with context (Constant-NotReinstated, Constant-Reinstated, and Varied) as a between-participants variable and measurement timing (Day 1 and Day 2) as a repeated measure. The result showed that only the main effect of measurement timing was significant (F (1, 58) = 132.9, p $<$ .001, ${\eta_{p}}^{2}$ = 0.70). Conversely, the main effect of context and the interaction were not significant (F (2, 58) = 1.09, p = .34, ${\eta_{p}}^{2}$ = 0.04; F (2, 58) = 0.37, p = .69, ${\eta_{p}}^{2}$ = 0.01).
@@ VI-A No reinstatement effect using IVE context
In this study, we investigated the reinstatement effect on all memory measures (RP1, acquisition, retention, and forgetting) and found no significant effects. Therefore, the hypotheses were not supported. Whereas the previous studies used 3DCG IVEs [9, 7, 6], this study is the first to use live-action 360-degree video for IVE. We assumed that increased realism and familiarity would improve the effectiveness of the environmental context, but the assumption was not supported.
@@
For RP1 and acquisition, the hypothesis stated that the number of correct responses obtained would be higher in the same context conditions (Constant-NotReinstated and Constant-Reinstated conditions), where tests were conducted in the same context as the initial study than in the different context condition (Varied condition), where tests were conducted in a new context. From Fig. 5, we observed no differences between the conditions, which were very different from those observed by Smith and Handy [5, 14] who showed the superiority of the same context condition to the different context condition by video/photo context manipulation in a similar procedure. Particularly in RP1, the number of correct answers was not significantly different between the same and different context conditions with the small effect size (same context: M = 3.71, SD = 2.52; different context: M = 4.05, SD = 3.01; Cohen’s d = 0.12).
@@
For retention, the hypothesis stated that the number of correct responses would be higher in the Constant-Reinstated condition, where participants performed the final test in the same context as Day 1, than in the Constant-NotReinstated condition, where the participants performed the final test in a context different from that of Day 1. Similarly, for forgetting, the hypothesis stated that forgetting would be lower in the Constant-Reinstated condition than in the Constant-NotReinstated condition. From Fig. 5 and Fig. 6, we observed no significant differences between the conditions, and the hypothesis was not supported. The small effect size suggests that the detection power may have been insufficient.
@@
The lack of effect may have been caused by cue overload [42], which is a common phenomenon observed in this field of study. Although the use of too many words eliminates the reinstatement effect, no such occurrence was analyzed in our experiment because Smith and Handy [14], who also used 20 Tagalog-English pairs of words, detected the effect. On the other hand, some procedural differences were observed between our study and that of Smith and Handy [14]. For example, Smith and Handy quickly switched the background photo contexts (one photo per word) in their study, while we used the same context for 20 items. The experimental protocol was modified because quick changes in IVE could confuse the participants and compromise the sense of presence, which may be vital for inducing context-dependent effects using VR. This difference in cue load level was not problematic in that it avoided the problem of cue overload but may have reduced the reinstatement effect [12]. However, Wälti et al. [6] already proved that even in a method like Smith and Handy’s, where the contexts are changed quickly per word, no reinstatement effect could be obtained in context manipulation by VR. Therefore, we do not consider that cue overload can explain the null results.
@@
Instead, the unique features of VR may have weakened the effect. From the discussion presented by Watson and Gaudl [8], we deduced that the congruence of the meta-cognitive context of “experiencing VR” may have counteracted the effects of visual and auditory context changes. This “experiencing VR” context was thought to be formed by wearing the HMD and high arousal during the VR experience. If this assumption is correct, the context cannot be changed irrespective of the extent of manipulation of audio-visual environmental information in VR. Similarly, the congruence of the meta-cognitive context of “being at the same university and undergoing the same experiment” was suggested to weaken the context-dependent effect when using the room as the environmental context [27]. However, despite the meta-cognitive context match, ways in which the effect can be stably detected have been suggested, for example, by using the background video context [12] or simultaneous manipulation of place, task, and social context [43]. Moreover, as the decontextualization effect was confirmed in the present study, the meta-cognitive contextual congruence of experiencing VR is suggested not to be fatal. Therefore, further research is expected to reveal conditions under which the reinstatement effect occurs using VR despite the meta-cognitive context match.
@@
One method to solve the meta-cognitive consistency problem could employ a further increase in presence. For example, increasing body tracking levels may be considered effective for enhancing presence [30]. As only three tracking points were used (the head and both hands) in this study, presence can be improved by attaching trackers to the waist and feet of the participants to achieve full-body tracking. Further, if the movements of the fingers can be faithfully reflected in the avatar by hand tracking, the presence can be improved. In this study, we did not observe a statistically significant impact of presence on reinstatement effects. However, if we continue our investigation with presence as the primary variable, we may observe an effect because both context-dependency of episodic memory and the sense of presence in IVE are dependent on hippocampal activity [13, 30].
@@ VI-B Decontextualization effect using IVE context
The second difference between our study and that of Watson and Gaudl [8] includes the type and number of IVEs used. Watson and Gaudl used two photorealistic 3DCG indoor environments, while we used seven 360-degree-video outdoor environments. Based on the encoding variability hypothesis, the more environmental contexts used, the more cues were available. Thus, the difference in the number of contexts might have improved the recall performance [44]. Additionally, the differences between 3DCG and 360-degree video, indoor and outdoor, can also be verified in the future.
Finally, Watson and Gaudl used doors for a contextual change, as doors have been found to act as boundaries for episodic memory [45]. Conversely, the participants of this study remained seated and only put on and off their HMD. Therefore, the possibility that putting on and off the HMD acted as an episodic boundary similar to a door or visually representing movement (e.g., virtual doors or portals) in the IVE may have improved the effect remains to be explored.
Although we did not observe a statistically significant effect of presence on the decontextualization effect, we are expecting the presence to be a possible parameter for improving the memory support effect observed in this study. According to the aforementioned discussion on the reinstatement effect, we can manipulate presence as an independent variable for validation in the future; for instance, by manipulating the tracking level.
Finally, the VR sickness measured by the SSQ did not differ significantly between conditions, as shown in Fig. 9. Therefore, the characteristic was not confounded for memory performance. VR sickness is one of the critical factors to be considered during practical VR-based applications. Therefore, this study is noteworthy in that we measured VR sickness for the first time in environmental context-dependent memory research using VR and showed relatively low severity.
@@ VI-C Limitations
Additionally, this study demonstrated a new method of foreign language learning as a clear application by employing the task of memorizing Japanese-Tagalog pairs. However, the generalizability of this research findings may go beyond language learning. For example, Smith and Handy (2014) showed decontextualization effects in a face/name pair memory task [5], and it is highly possible that our findings can be applied to various areas as long as classified in the field of episodic memory. On the other hand, the results are unlikely to be applicable to memories that cannot be described verbally, such as memories of motor skills. We employed language learning to be consistent with the use of word lists as the material to be memorized in most previous studies and also because Smith and Handy (2016) [14], who found decontextualization effects, employed Tagalog-English pair learning. Future studies are expected to conduct validation in various situational settings to clarify the generalizability of the results.
@@ VI-D Future Work
In this experiment, it was verified that while the reinstatement effect using VR is unlikely to occur, the decontextualization effect does occur. As the decontextualization effect using VR was revealed for the first time in this study, replication and follow-up experiments should be conducted to elucidate the conditions under which the effect can be enhanced. For example, the relationship between presence and context-dependent effect is an essential topic for future work.
@@ VII Conclusion
This study tested whether context-dependent forgetting can be eliminated by inducing reinstatement and decontextualization effects using 360-degree video-based virtual environments. For the reinstatement effect, complex results were obtained, suggesting the need for further elaborate research. On the other hand, for the decontextualization effect, this study was the first to validate the effect using VR and revealed its statistically significant effect with a large effect size. While previous studies have assumed that VR could naturally reproduce the context-dependent effects observed in the real world, most of them could not prove the assumption. Given this context, obtaining a clear decontextualization effect in this study sheds light on the possibility of reproducing context-dependent effects using VR.